Assume that the spot exchange rate of the British pound is $1.73/£. How will this spot rate adjust according to PPP if the United Kingdom experiences an inflation rate of 7 percent while the United States experiences an inflation rate of 2 percent?
Answer #1

Lets say an item is sold in US for 1.73 dollars

and the same item is sold in UK for 1 pound

After 1 year,
In US price=1.73*1.02 dollar
In UK price=1*1.07 pound

Hence, spot rate 1 year later=1.73*1.02/1.07=1.649158879 $/pound